OpenStudy (meszpilka):

using angle sum identities, how does cos 165 = sq root (2 + sq root 3)/2?

OpenStudy (anikhalder):

cos(165)=cos(120+45) =cos120cos45-sin120sin45 =cos(90+30)cos45-sin(90+30)sin45 Now, cos(90+30)=-sin30 sin(90+30)=cos30 And values of cos45 and sin45 are equal so we can write cos45=sin45 and take a common in the equation say..cos45 =cos45(-sin30-cos30) =-cos45(sin30+cos30) Putting in values we get the value of the exp to be=\[-\frac{ 1 }{ \sqrt{2}}(\frac{ 1 }{ 2}+\frac{ \sqrt{3} }{ 2})\] Nowhere near the proof..have I made any mistake?